QUESTION 5 Past records indicate that each of the 5 laser computer printers at the U.S. Department of Energy needs repair aftFind the average waiting time.

Average number waiting.

Average number waiting.

Average number being serviced.
Answer #1

Given that,

Number of printers, M = 5

Arrival rate, \lambda = 1/20 = 0.05 per hour

Service rate, \mu = 1/2 = 0.5 per hour

P0 = 0.5640

-----------------------------------------

Average utilization, \rho = 1 - P0 = 1-0.5640 = 0.4360

Average number waiting, Lq = N - (\lambda+\mu)\rho/\lambda = 5 - (0.05+0.5)*0.436/0.05 = 0.204

Average number in the system, L = Lq+\rho = 0.204+0.436 = 0.64

Average waiting time, Wq = Lq/((N-L)*\lambda) = 0.204/((5-0.64)*0.05) = 0.93 hour

Average number being serviced = \rho = 0.436